OpenStudy (anonymous):

\[\sqrt{1-\cos \alpha/1+\cos \alpha}=\csc \alpha-\cot \alpha\]

OpenStudy (anonymous):

identities

OpenStudy (jamesj):

Square both sides and try.

OpenStudy (anonymous):

with this kind of problem you're supposed to work on 1 side only

OpenStudy (jamesj):

With this sort of problem you're not going to get anywhere with the square root there, so try and get rid of it first and see what you get.

OpenStudy (anonymous):

Doesn't actually seem to be an identity? See: http://www.wolframalpha.com/input/?i=Plot [{Csc[x]-Cot[x]%2C+%28%281-Cos[x]%29%2F%281%2BCos[x]%29%29^.5}%2C+{x%2C0%2C10*Pi}] Squaring both sides would make it look like one when it is not.

OpenStudy (anonymous):

For example try \[\alpha=3\pi/2\] Csc[3*Pi/2] - Cot[3*Pi/2] = -1 Sqrt[(1 - Cos[3*Pi/2])/(1 + Cos[3*Pi/2])] = 1